OpenStudy (anonymous):

Equation of a circle given a center of (-5, 2) and a radius of 5.

hartnn (hartnn):

Equation of circle with center (h,k) and radius 'r' will be \(\large (x-h)^2+(y-k)^2 = r^2\) for your question, h= -5, k = 2, r = 5 just plug them in! and simplify :)

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^